# Trapping of Single Atoms in Metasurface Optical Tweezer Arrays

### v2

@@ Abstract
Optical tweezer arrays have emerged as a key experimental platform for quantum computation, quantum simulation, and quantum metrology, enabling unprecedented levels of control over single atoms and molecules. Existing methods to generate tweezer arrays mostly rely on active beam-shaping devices, such as acousto-optic deflectors or liquid-crystal spatial light modulators. However, these approaches have fundamental limitations in array geometry, size, and scalability. Here we demonstrate the trapping of single atoms in optical tweezer arrays generated via holographic metasurfaces. We realize two-dimensional arrays with more than 250 tweezer traps, arranged in arbitrary geometries with trap spacings as small as 1.5 µm. The arrays have a high uniformity in terms of trap depth, trap frequency, and positional accuracy, rivaling or exceeding existing approaches. Owing to sub-micrometer pixel sizes and high pixel densities, holographic metasurfaces open a path towards optical tweezer arrays with $>100,000$ traps, facilitating tweezer-array based quantum applications that require large system sizes.
@@
Fig. 1: Metasurface optical tweezer arrays and integration with ultracold strontium atoms. a, Illustration of the working principle of a holographic metasurface. The metasurface imprints a phase pattern on an incident collimated Gaussian laser beam and produces a two-dimensional array of tight tweezer traps at the focal plane. In this work, the focal length of the metasurfaces is 0.7 mm and the effective NA is $>0.6$. b, (left) Photograph of a substrate holding 6 metasurfaces, each with a distinct tweezer array pattern. The substrate has a size of 20 mm $\times$ 20 mm. The metasurfaces have a diameter of 1.16 mm. (middle) Scanning electron microscope image of a portion of a metasurface. (right) The meta-atoms have a unit cell size of 290 nm, are 100–190 nm wide, and 750 nm tall. They are chosen from a predetermined library to introduce a wide range of phase delays to the incoming beam without modifying its amplitude. c, Level diagram of 88Sr showing the optical transitions relevant to this work. Atoms are cooled and imaged on the broad transition at 461 nm in conjunction with the repumpers at 679 nm and 707 nm. Narrow-line cooling on the intercombination line at 689 nm (linewidth 7.5 kHz) creates Sr samples at microkelvin temperatures. d, Schematic of the setup to trap and image atoms. The tweezer laser operates at 520 nm. Its intensity is controlled with an acousto-optic modulator (AOM) before illuminating the metasurface. The metasurface generates and focuses the tweezer array, which is collimated by a high-NA microscope objective, relayed via a 1:1 telescope, and focused down with a high-NA objective into the ultra-high vacuum glass cell to trap atoms. During imaging, fluorescence photons of the atoms are detected with a single-photon sensitive camera.
@@
The ability to generate high-quality optical tweezer arrays is a central requirement for many experiments. A tweezer array consists of numerous tightly focused laser beams, each constituting a trap for a single particle. Key criteria include high flexibility in array geometry, trap uniformity, and scalability. In addition, compactness, robustness, and high optical efficiency are desirable, especially with the prospect of deploying tweezer-based quantum devices outside of controlled laboratory environments [22, 23, 24]. Currently, optical tweezer arrays are mostly generated via active beam-shaping devices, such as acousto-optical deflectors (AODs) [25, 26], liquid crystal spatial light modulators (SLMs) [27, 28], or digital micromirror devices (DMDs) [29]. These devices require complex control electronics and projection optics with high numerical aperture (NA) to relay the tweezer arrays onto ultracold atoms and molecules. Technical complexity and fundamental limitations constrain array sizes to $<10,000$ traps, which has started to impose a limit on the quantum applications that can be pursued. Alternative techniques, such as amplitude masks [30] and microlens arrays [31], have been explored, but limited beam-shaping capabilities make it challenging to achieve highly uniform arrays.
@@
In recent work [32], we proposed holographic metasurfaces as a new approach to generate versatile and scalable tweezer arrays. Metasurfaces are flat optical devices comprised of sub-micrometer pixels that can imprint an arbitrary phase mask onto an incident laser beam [33, 34, 35], both generating and focusing an optical tweezer array. Metasurfaces feature high power handling capabilities [36], diffraction-limited focusing [37], and comprehensive polarization control [38, 39]. A recent experiment demonstrated single atom trapping in a $3\times 3$ tweezer array that was generated with an AOD and focused down by a metasurface lens [40]; however, a demonstration of atomic tweezer arrays that utilize the full beam-shaping capabilities of metasurfaces – array generation and focusing – has so far been elusive.
@@
In this work, we demonstrate the trapping of single strontium (Sr) atoms in metasurface optical tweezer arrays. Using laser light at a wavelength of 520 nm, we realize two-dimensional arrays with arbitrary geometries, both periodic and non-periodic, with more than 250 traps and trap spacings as small as 1.5 µm. We find the trap uniformity to be comparable to state-of-the-art techniques and demonstrate single-atom preparation and detection with high fidelity. The holographic metasurfaces in this work have sub-micrometer and subwavelength pixel sizes, offering a realistic path towards array sizes beyond 100,000 traps, and thus overcoming fundamental limitations of existing beam-shaping approaches.
@@
Metasurfaces have emerged as a powerful platform for the manipulation of optical waves [41, 42]. They enable holographic control of incident light fields by manipulating the amplitude and phase of an optical wavefront in the device plane. Metasurfaces are comprised of a matrix of nano-etched meta-atoms that serve as control pixels; the meta-atoms are selected from a library – a collection of scatterers with different shapes and sizes – and carefully positioned to engineer the optical wavefront that leads to the desired pattern in the imaging plane. Each meta-atom is a few hundred nanometers wide and tall, smaller than the wavelength of the light they manipulate. Metasurfaces can be fabricated on the millimeter to centimeter scale [43, 44] with well over $10^{6}$ pixels. Because of the subwavelength size of the meta-atoms, metasurfaces can achieve high spatial resolution and forward-scattering efficiency.
@@
We use transmitting metasurfaces for visible light at 520 nm (Fig. 1a). These metasurfaces are phase-only modulating masks designed using a Gerchberg-Saxton algorithm-based optimization approach (see Methods for details). The optimized metasurface encodes a phase pattern that simultaneously generates and focuses a tweezer array. Multiple metasurfaces are placed on a single substrate (Fig. 1b), allowing for easy switching between distinct designs and their associated tweezer arrays by translation of the substrate. The metasurfaces are fabricated using standard cleanroom protocols from thin films of silicon-rich silicon nitride (SRN). This material is chosen for its high optical refractive index and ease of manufacturing (see Methods for additional considerations). To minimize alignment sensitivity and mitigate diffraction effects of the Gaussian input beam at the metasurface boundaries, the metasurface is manufactured with a circular footprint and assumes the incidence of a plane wave onto the device. The metasurfaces have a diameter of 1.16 mm and 2.32 mm. They can handle optical intensities up to 25 W/mm2 without active cooling, have a diffraction efficiency of $\sim 60\%$, and an effective NA of $>0.6$. Additional details on the design and optimization of the metasurface hologram, meta-atom library, cleanroom fabrication procedure, and optical properties are provided in the Methods.
@@
Fig. 2: Fluorescence images of strontium atoms trapped in metasurface-generated optical tweezer arrays. Each image is an average of 100 individual images without parity projection. a, Square lattice with 256 traps, average spacing of 4 µm. b, Quasicrystal (Penrose tiling) with 225 traps, average spacing of 4 µm. c, Statue of Liberty pattern with 183 traps, average spacing of 3 µm. d, Necklace pattern with 16 traps, average spacing of 1.45 µm.
@@
The optical tweezer arrays are projected into the glass cell of an ultrahigh-vacuum chamber (see Methods for details). A schematic of the projection system is shown in Fig. 1d. Before illuminating the metasurface, a laser at 520 nm passes through an acousto-optic modulator (AOM) for fast switching and trap depth control. The substrate with the metasurfaces is mounted on a two-axis translation stage, allowing for rapid switching between different array geometries with minimal realignment. The tweezer array, generated by the metasurface at the focal plane, is converted into the optical momentum space by a microscope lens ($\text{NA}=0.6$), relayed through a 1:1 telescope, and converted back into the tweezer array in the glass cell by an objective lens ($\text{NA}=0.5$). Tweezer generation with an AOD, liquid crystal SLM, or DMD typically requires demagnification optics with large-diameter lenses that are prone to aberrations. Due to the high NA of the metasurface, demagnification is not necessary. In principle, metasurfaces can directly trap atoms at their focal plane (e.g., by placing them inside or near the vacuum chamber) without the need for additional relay optics.
@@
We load atoms into the metasurface optical tweezer array from an ultracold cloud of 88Sr. The strontium atoms are cooled to microkelvin temperatures using standard techniques that leverage strontium’s unique level structure (Fig. 1c). Subsequently, the trapped atoms are detected via fluorescence imaging (see Methods). Figure 2 shows fluorescence images of atoms in different metasurface-generated tweezer arrays, including a periodic $16\times 16$ pattern, a quasicrystal pattern (Penrose tiling), a fully arbitrary pattern, and a necklace pattern with close tweezer spacing on the micrometer-scale.
Single atom trapping and imaging
@@
We demonstrate single-atom trapping and detection in a $4\times 4$ metasurface array (see Fig. 3). The steps to achieve this are highly sensitive to the quality and uniformity of the optical tweezer potential. The initial loading of the array is statistical; each trap is occupied by at least one atom, but the precise number of atoms is random. In the next step, we perform parity projection: traps that initially have an odd (even) number of atoms are turned into sites with one (no) atom. This is achieved via photoassociation into an electronically excited Sr2 molecular state, close to the 689 nm atomic resonance, which induces pairwise atom loss [45]. After parity projection, $49(3)\%$ of the traps contain an atom, as shown in Fig. 3b.
@@
Fig. 3: Single atom preparation and detection in a $4\times 4$ metasurface array. a, Average of 100 fluorescence images after parity projection. The uniformity indicates that all traps have an approximately equal chance of being filled with an atom. b, (left) Individual fluorescence image. Trap locations are indicated by dashed boxes. (right) Histogram of the number of occupied sites after parity projection; the mean occupancy of a trap, marked by the dashed line, is $49(3)\%$. c, Histogram of photon counts in the trap locations of the $4\times 4$ array, as marked in b. 500 repetitions of the experiment are averaged. The data allows for the distinction between sites with one and zero atoms with high fidelity. The dashed line marks the threshold value, determined as discussed in Methods.
@@
To determine the occupation in the array, we perform fluorescence imaging on the 461 nm transition, while keeping the atoms trapped. Fluorescence photons are collected with a low-noise camera and the photon number in the trap locations (marked in Fig. 3b, left panel) is evaluated. To maximize the number of photons scattered per atom, we simultaneously cool via repulsive Sisyphus cooling [46] on the 689 nm, $m_{J}=\pm 1$ transition, counteracting the recoil heating from repeated photon scattering on the 461 nm transition. Figure 3c shows a histogram of the detected photon numbers in the trap locations. The histogram shows two peaks: one peak centered on zero photons, corresponding to zero atoms, and a second peak centered on $\sim 8$ photons, corresponding to the presence of a single atom. The absence of photon counts above the single-atom peak indicates the high efficiency of parity projection. The presence of photon counts between the zero- and single-atom peaks results from loss of Sr atoms during imaging in 520 nm traps. Similar observations were reported in Ref. [47]. We attribute this loss to an ionization process out of the ${}^{3}\text{P}_{1}$ state, which is populated during Sisyphus cooling. This will be analyzed in further detail in future work. For alternative trapping wavelengths, for example 813 nm [48], such losses are known to be absent. Using a model-free approach that involves the recording of two fluorescence images in short sequence [49], we determine the imaging fidelity to be $>99.8(5)\%$ (see Methods).
Characterization of array quality
@@
Next, we characterize the uniformity of a $16\times 16$ array. We use the trapped atoms as highly sensitive probes to measure the depth, frequency, and position of each trap. A high uniformity in trap depth and frequency ensures that the light shift and on-site vibrational modes are constant across the array. These characteristics are important for sideband cooling, atomic state preparation, and spectrally sensitive operations, such as atomic quantum gates and interrogation of clock states. A high accuracy of trap positions is desirable for the precise control of atom-atom interactions, for example, in Rydberg arrays and correlated atom-photon interactions.
We characterize the uniformity of trap depth by measuring the resonance frequency of the narrow Sr intercombination transition at 689 nm in each tweezer. The resonance is shifted away from the free-space resonance due to a differential light shift between the 1S0 and the 3P1 states in the 520 nm trapping field. While keeping the overall laser intensity fixed, the 689 nm resonance is measured in each trap via loss spectroscopy. From this, we determine the local light shift, which, with the knowledge of the polarizability of 1S0 and 3P1, is converted into a trap depth. As shown in Fig. 4a, we find a mean trap depth of $0.52$ mK with a standard deviation of $7.5\%$ across the array.
@@
The trap frequencies of individual tweezers are measured via parametric driving, both in the radial and axial directions. The intensity of the tweezer light is modulated sinusoidally with an amplitude of $5\%$ for 30 ms while the frequency of the modulation is varied. A parametric resonance is observed via atom loss at twice the trap frequency. Averaged over the array, we measure a mean radial (axial) trap frequency of $105$ kHz ($14$ kHz) with a standard deviation of $5\%$ $(8\%)$. Figure 4b shows the data for the radial trap frequency. The data for the axial trap frequency is shown in Extended Data Fig. 7.
To characterize the spatial accuracy, we compare the measured trap locations with the target trap positions that were used for the design of the metasurface. Fitting each atom’s fluorescence signal to a 2D Gaussian, we extract the tweezer locations. The positional inaccuracy between intended and measured locations is defined as $\Delta r=\sqrt{\Delta x^{2}+\Delta y^{2}}$, where $\Delta x$ ($\Delta y$) is the deviation in the $x$ ($y$) direction. We find a mean deviation of 60 nm, as shown in Fig. 4c. Compared with the trap spacing of 4 µm, this corresponds to a positional inaccuracy on the $1.5\%$ level. The deviation is comparable to the extent of the ground-state vibrational wavefunctions of Sr of about 30 nm and 90 nm in the radial and axial direction, respectively.
The uniformity of trap depth, frequency, and position of the metasurface approach rivals or exceeds the performance of existing techniques. For example, arrays generated with liquid crystal SLMs often have trap depth fluctuations of 10% to 20% [50, 28, 51], while methods to further reduce such fluctuations via feedback are under active research [52, 53]. For the metasurface arrays, we attribute the remaining non-uniformity to imperfections in the optical relay optics and to fabrication errors of meta-atoms, both of which have considerable room for improvement. Prior to transmission through the relay optics, we measure an intensity non-uniformity at the focal plane of the metasurface as low as 4% (see Methods).
@@
Fig. 5: Performance of pixel-based beam shaping devices. a, Approximation of the phase profile of a spherical lens with infinitely small pixel size ($d\ll\lambda$), intermediate pixel size ($d\sim\lambda$), and large pixel size ($d\gg\lambda$). For larger pixel sizes, the reproduction of steeper phase gradients $\partial\phi/\partial x$ is limited. This limits the usable diameter of the device and reduces the effective NA. b, Effective NA of a lens generated with a pixel-based device as a function of pixel density $1/d$ for common trapping wavelengths $\lambda$. The dashed vertical lines indicate the approximate separation between the regimes $d\ll\lambda$, $d\sim\lambda$, and $d\gg\lambda$. The yellow star indicates the pixel density for the metasurfaces used in this work. c, Focusing capabilities of pixel-based beam shaping devices. For a laser wavelength of 520 nm, a fixed device resolution of $300\times 300$ pixels, and varying pixel density $1/d$, an optimized $3\times 3$ square array with 5 µm spacing is generated (insets) (further details in Methods). Data points show the tightness of the traps, measured as the FWHM. Error bars show the standard deviation across the array. The pixel densities range from state-of-the-art liquid crystal SLMs ($d=4$ µm) to the holographic metasurface used in this work (yellow star, $d=290$ nm). The dashed line shows a fit of the effective NA model (further details in Methods). d, Simulation of the uniformity of trap intensity as a function of the number of tweezer traps for device resolutions ranging from $000\times 1,$ 000 × 1 , 000 (light green) to $000\times 16,$ 000 × 16 , 000 (dark blue) pixels (square-shaped device). Uniformity is defined as 100% minus the standard deviation of the trap intensity across the array (in %). The simulation assumes a pixel size of $d=290$ nm (further details in Methods).
@@
In this section, we discuss the scalability of metasurface optical tweezer arrays and show how their performance compares to other pixel-based beam shaping devices, such as DMDs and liquid crystal SLMs. We find that pixel size and number have a significant impact on the ability to create large and uniform tweezer arrays.
@@
To illustrate this point, we first consider reproducing a simple lens with a pixel-based device, as shown in Fig. 5a. The pixel size is denoted by $d$, the pixel density is $1/d$, and the wavelength of the incident laser beam is denoted by $\lambda$. We first convert a spherical lens to a Fresnel lens by taking the phase profile $\phi(x)$ of the former, modulo $2\pi$. When the phase profile of the Fresnel lens is approximated with a pixel-based device, the finite sampling constrains the steepness of phase gradients that can be reproduced, $\partial\phi/\partial x\lesssim\pi/2d$. This limits the ability to reproduce steep phase gradients at the edge of the lens. A larger pixel size will reduce the attainable phase gradient, reducing the usable diameter of the device, and effectively reducing the NA that can be achieved. Based on this argument, we derive an approximate analytical expression for the effective numerical aperture attainable for a given pixel size, $\mathrm{NA}=1/\sqrt{(1+(4d/\lambda)^{2})}$ (see Methods for details). In Fig. 5b, we plot the effective NA for a broad range of pixel densities and several common tweezer wavelengths in the visible and near-infrared, illustrating that a smaller pixel size relative to the laser wavelength can lead to a dramatically increased effective NA. For a pixel size of several µm (typical for DMDs and liquid crystal SLMs), in the regime $d\gg\lambda$, the NA is limited to below 0.05. For subwavelength pixels, in the regime $d\lesssim\lambda$, NAs of 0.5 and higher can be reached.

### v3

@@ Abstract
Optical tweezer arrays have emerged as a key experimental platform for quantum computation, quantum simulation, and quantum metrology, enabling unprecedented levels of control over single atoms and molecules. However, existing tweezer platforms have fundamental limitations in array geometry, size, and scalability. Here we demonstrate the trapping of single strontium atoms in optical tweezer arrays generated via holographic metasurfaces. We realize two dimensional arrays with more than 1000 trapped atoms, arranged in arbitrary geometries with trap spacings as small as 1.5 µm. The arrays have a high uniformity in terms of trap depth, trap frequency, and positional accuracy, rivaling or surpassing existing approaches. This is enabled by highly efficient holographic metasurfaces fabricated from high-refractive index materials, silicon-rich silicon nitride and titanium dioxide. Leveraging sub-micrometer pixel sizes and high pixel densities, our platform allows scaling far beyond current capabilities. As a demonstration, we realize an optical tweezer array with 360,000 traps. These advances will facilitate tweezer-array based quantum applications that require large system sizes.
@@
Fig. 1: Metasurface optical tweezer arrays and integration with ultracold strontium atoms. a, Illustration of the working principle of a holographic metasurface. The metasurface imprints a phase pattern on an incident collimated Gaussian laser beam and produces a two-dimensional array of tight tweezer traps at the focal plane. In this work, the focal length of the metasurfaces is 0.7 mm and the effective NA is >0.6>0.6. b, (left) Photograph of a substrate holding 6 metasurfaces, each with a distinct tweezer array pattern. The substrate has a size of 20 mm $\times$ 20 mm. The metasurfaces have a diameter of 1.16 mm. (middle) Scanning electron microscope image of a portion of a metasurface. (right) The meta-atoms have a unit cell size of 290 nm, are 100–190 nm wide, and 750 nm tall. They are chosen from a predetermined library to introduce a wide range of phase delays to the incoming beam without modifying its amplitude. c, Level diagram of 88Sr showing the optical transitions relevant to this work. Atoms are cooled and imaged on the broad transition at 461 nm in conjunction with the repumpers at 679 nm and 707 nm. Narrow-line cooling on the intercombination line at 689 nm (linewidth 7.5 kHz) creates Sr samples at microkelvin temperatures. d, Schematic of the setup to trap and image atoms. The tweezer laser operates at 520 nm. Its intensity is controlled with an acousto-optic modulator (AOM) before illuminating the metasurface. The metasurface generates and focuses the tweezer array, which is collimated by a high-NA microscope objective, relayed via a 1:1 telescope, and focused down with a high-NA objective into the ultra-high vacuum glass cell to trap atoms. During imaging, fluorescence photons of the atoms are detected with a single-photon sensitive camera.
@@
The ability to generate high-quality optical tweezer arrays is a central requirement for many experiments. A tweezer array consists of numerous tightly focused laser beams, each constituting a trap for a single particle. Key criteria for the platform include high flexibility in array geometry, trap uniformity, and scalability. In addition, compactness, robustness, and high optical efficiency are desirable, especially with the prospect of deploying tweezer-based quantum devices outside of controlled laboratory environments [22, 23, 24]. Currently, optical tweezer arrays are mostly generated via active beam-shaping devices, such as acousto-optical deflectors (AODs) [25, 26], liquid crystal-spatial light modulators (SLMs) [27, 28], or digital micromirror devices (DMDs) [29]. These devices require complex control electronics and projection optics with high numerical aperture (NA) to relay the tweezer arrays onto ultracold atoms and molecules. Technical complexity and fundamental limitations constrain array sizes to $\sim 10,000$ traps [30], which has started to impose a limit on the quantum applications that can be pursued. Alternative techniques, such as amplitude masks [31] and microlens arrays [32], have been explored, but limited beam-shaping capabilities make it challenging to achieve highly uniform arrays.
@@
In recent work [33], we proposed holographic metasurfaces as a new approach to generate versatile and scalable tweezer arrays. Metasurfaces are flat optical devices comprised of sub-micrometer pixels that can imprint an arbitrary phase mask onto an incident laser beam [34, 35, 36], both generating and focusing an optical tweezer array. Metasurfaces feature high power-handling capabilities [37], diffraction-limited focusing [38, 39], and comprehensive polarization control [40, 41]. A recent experiment demonstrated single atom trapping in a $3\times 3$ tweezer array that was generated with an AOD and then focused down by a metasurface lens [42]; however, a demonstration of atomic tweezer arrays that utilize the full beam-shaping capabilities of metasurfaces – integrating array generation and focusing into one device – has so far been elusive.
@@
In this work, we demonstrate the trapping of single strontium (Sr) atoms in metasurface optical tweezer arrays. Using laser light at a wavelength of 520 nm, we realize two-dimensional atomic arrays with arbitrary geometries, both periodic and non-periodic, with more than 1000 trapped atoms and trap spacings as small as 1.5 µm. We find the trap uniformity to be comparable to state-of-the-art techniques and demonstrate single-atom preparation and detection with high fidelity. We discuss how large-area holographic metasurfaces with subwavelength pixel sizes offer a realistic path towards tweezer arrays with >100,000 atoms. Finally, we experimentally demonstrate a highly uniform optical tweezer array with 360,000 traps, exceeding the current state-of-the-art by two orders of magnitude.
@@
Metasurfaces have emerged as a powerful platform for the manipulation of optical waves [43, 44]. They enable holographic control of incident light fields by manipulating the amplitude and phase of an optical wavefront in the metasurface plane. Our metasurfaces are composed of nanofabricated meta-atoms, each a dielectric nanopillar, a few hundred nanometers in width and height, smaller than the wavelength of the light they manipulate. Meta-atoms with different shapes and sizes are positioned in a two-dimensional (2D) grid with subwavelength spacing to engineer the optical wavefront and generate the desired optical intensity pattern in the imaging plane. Metasurfaces can be fabricated on the millimeter to centimeter scale [45, 46] with well over $10^{6}$ pixels. Due to the subwavelength size and spacing of the meta-atoms, metasurfaces can achieve high optical efficiency and high precision of the generated tweezer pattern.
@@
We use transmitting metasurfaces for visible light at 520 nm (Fig. 1a). These metasurfaces are phase-only modulating masks designed using a Gerchberg-Saxton algorithm-based optimization approach (see Methods for details). The optimized metasurface encodes a phase pattern that simultaneously generates and focuses a tweezer array. Multiple metasurfaces are placed on a single substrate (Fig. 1b), allowing for easy switching between distinct tweezer arrays by translation of the substrate. The metasurfaces are implemented on two complementary dielectric platforms: silicon-rich silicon nitride (SRN) - allowing for fast CMOS-compatible fabrication - and titanium dioxide ($\text{TiO}_{2}$) - featuring superior power handling and compatibility with shorter optical wavelengths. The metasurfaces are designed assuming the incidence of a flat phase front and have a circular footprint to reduce diffraction effects of round input beams at their boundaries. The metasurfaces have diameters ranging from 1.2 mm to 3.5 mm. They can handle optical intensities of at least 25 W/mm2 (SRN) or 2,000 W/mm2 ($\text{TiO}_{2}$) without active cooling, have a diffraction efficiency of $\sim 60\%$, and an effective NA of >0.6>0.6. Additional details on the design and optimization of the metasurface hologram, meta-atom library, cleanroom fabrication procedure, and a comparison of holographic metasurfaces with AODs and SLMs are provided in the Methods.
@@
Fig. 2: Fluorescence images of strontium atoms trapped in metasurface-generated optical tweezer arrays. Each image is an average of 100 individual images without parity projection. a, Statue of Liberty pattern with 183 traps, average spacing of 3 µm. b, Quasicrystal (Penrose tiling) with 225 traps, average spacing of 4 µm. c, Square lattice with 1024 traps, average spacing of 2.5 µm. d, Necklace pattern with 16 traps, average spacing of 1.45 µm.
@@
The optical tweezer arrays are projected into the glass cell of an ultrahigh-vacuum chamber (see Methods for details). A schematic of the projection system is shown in Fig. 1d. Before illuminating the metasurface, a laser at 520 nm passes through an acousto-optic modulator (AOM) for fast switching and trap depth control. The substrate with the metasurfaces is mounted on a two-axis translation stage, allowing for rapid switching between different array geometries with minimal realignment. The tweezer array, generated by the metasurface at the focal plane, is converted into the optical momentum space by a microscope lens ($\text{NA}=0.6$), relayed through a 1:1 telescope, and converted back into the tweezer array in the glass cell by an objective lens ($\text{NA}=0.5$). Tweezer generation with AODs, SLMs, or DMDs typically requires demagnification optics with large-diameter lenses that are prone to aberrations and impose power-handling and field-of-view limitations; in contrast, for our metasurfaces demagnification optics are not necessary due to their intrinsically high NA. In principle, metasurfaces can directly trap atoms at their focal plane (e.g., by placing them inside or near the vacuum chamber) without the need for additional relay optics.
@@
We load atoms into the metasurface optical tweezer array from an ultracold cloud of 88Sr. The atoms are cooled to microkelvin temperatures using standard techniques that leverage strontium’s unique level structure (Fig. 1c). Subsequently, the trapped atoms are detected via fluorescence imaging (see Methods). Figure 2 shows fluorescence images of atoms in different metasurface-generated tweezer arrays including a fully arbitrary pattern (Statue of Liberty), a quasicrystal pattern (Penrose tiling), a periodic $32\times 32$ square lattice pattern, and a necklace pattern with close tweezer spacings on the micrometer-scale. The array sizes are limited to a few hundred trapped atoms solely by the available tweezer laser power of about 1 W in our current setup.
Single atom trapping and characterization
@@
We demonstrate single-atom trapping and detection in a $16\times 16$ metasurface array. The steps to achieve this are highly sensitive to the quality and uniformity of the optical tweezer potential. The initial loading of the array is statistical; each trap is occupied by at least one atom, but the precise number of atoms is random (see Fig. 3a). In the next step, we perform parity projection: traps that initially have an odd (even) number of atoms are turned into sites with one (no) atom. This is achieved via photoassociation into an electronically excited Sr2 molecular state, close to the 689 nm atomic resonance, which induces pairwise atom loss [47]. After parity projection, 41(4)% of the traps contain an atom, corresponding to >100>100 single atoms in the array, as shown in Fig. 3b.
@@
Fig. 3: Single atom preparation and detection in a $\mathbf{16\times 16}$ metasurface array. a, Individual fluorescence image. Trap locations are indicated by dashed boxes. b, Histogram of the number of occupied sites after parity projection; the mean occupancy of the array is 106(11), as marked by the dashed line, corresponding to a mean trap occupancy of 41(4)%. c, Histogram of photon counts for a typical trap in the $16\times 16$ array, as marked in a. 500 repetitions of the experiment are averaged. The data allows for the distinction between one and zero atoms with high fidelity. The dashed line marks the threshold value, determined as discussed in Methods.
@@
To determine the occupation in the array, we perform fluorescence imaging on the 461 nm transition, while keeping the atoms trapped. Fluorescence photons are collected with a low-noise camera and the photon number in the trap locations is evaluated. To maximize the number of photons scattered per atom, we simultaneously cool via repulsive Sisyphus cooling [48] on the 689 nm, $m_{J}=\pm 1$ transition, counteracting the recoil heating from repeated photon scattering on the 461 nm transition. Figure 3c shows a histogram of the detected photon numbers for a typical trap in the array. The histogram shows two peaks: one peak centered on zero photons, corresponding to zero atoms, and a second peak centered on $\sim 4.5$ photons, corresponding to a single atom. The absence of photon counts above the single-atom peak indicates the high efficiency of parity projection. The presence of photon counts between the zero- and single-atom peaks results from loss of Sr atoms during imaging in the 520-nm traps. Similar observations were reported in Ref. [49]. We attribute this loss to an ionization process out of the $^{3}\text{P}_{1}$ state, which is populated during Sisyphus cooling. This mechanism will be analyzed in further detail in future work. For alternative trapping wavelengths, for example 813 nm [50], such losses are known to be absent. We determine the imaging fidelity to be $95(3)\%$. When using a smaller $4\times 4$ array, not limited by tweezer laser power, we observe a filling fraction of $49(3)\%$ and an imaging fidelity of $99.8(5)\%$ (see Methods).
@@
Next, we characterize the uniformity of the $16\times 16$ array. A high uniformity in trap depth and frequency ensures that the light shift and on-site vibrational modes are constant across the array; a high accuracy of trap positions is desirable for the precise control of atom-atom interactions. We use the trapped atoms as highly sensitive probes to measure the depth, frequency, and position of each trap. Additional details on the measurements are provided in the Methods. Fig. 4 shows the results of our characterization. For the trap depth and the radial (axial) trap frequencies we find a standard deviation of $7.5\%$ and $5\%$ $(8\%)$, respectively, across the array. The positional inaccuracy is on the $1.5\%$ level compared to the trap spacing of 4 µm, similar to the extent of the in-trap vibrational wavefunction of Sr.
The data shows that the uniformity of our metasurface arrays rivals or surpasses the performance of existing techniques. For example, arrays generated with liquid crystal SLMs often have trap depth fluctuations of 10% [51, 30]. Methods to reduce such fluctuations via feedback are under active research [28, 52, 53], but are fundamentally limited for large-scale arrays as discussed in the next section. For the metasurface arrays, we attribute the remaining non-uniformity to imperfections in the relay optics and to fabrication errors of meta-atoms, both of which have considerable room for improvement. Prior to transmission through the relay optics, we measure an intensity non-uniformity at the focal plane of the metasurface as low as 4% (see Methods).
@@
Fig. 5: Performance of pixel-based beam shaping devices. a, Approximation of the phase profile of a spherical lens with infinitely small pixel size ($d\ll\lambda$), intermediate pixel size ($d\sim\lambda$), and large pixel size ($d\gg\lambda$). For larger pixel sizes, the reproduction of steeper phase gradients $\partial\phi/\partial x$ is limited. This limits the usable diameter of the device and reduces the effective NA. b, Effective NA of a lens generated with a pixel-based device as a function of pixel density $1/d$ for common trapping wavelengths $\lambda$. The dashed vertical lines indicate the approximate separation between the regimes $d\ll\lambda$, $d\sim\lambda$, and $d\gg\lambda$. The yellow star indicates the pixel density for the metasurfaces used in this work. c, Focusing capabilities of pixel-based beam shaping devices. For a laser wavelength of 520 nm, a fixed device resolution of $300\times 300$ pixels, and varying pixel density $1/d$, an optimized $3\times 3$ square array with 5 µm spacing is generated (insets) (further details in Methods). Data points show the tightness of the traps, measured as the FWHM. Error bars show the standard deviation across the array. The pixel densities range from state-of-the-art liquid crystal SLMs ($d=4$ µm) to the holographic metasurface used in this work (yellow star, $d=290$ nm). The dashed line shows a fit of the effective NA model (further details in Methods). d, Simulation of the uniformity of trap intensity as a function of the number of tweezer traps for device pixel counts ranging from $1,000\times 1,000$ (light green) to $16,000\times 16,000$ (dark blue) pixels (square-shaped device). Uniformity is defined as 100% minus the standard deviation of the trap intensity across the array (in %). The simulation assumes a pixel size of $d=290$ nm (further details in Methods).
Fig. 6: Realization and characterization of a $\mathbf{600\times 600}$ tweezer array. a, (top) Setup for optical characterization including a high NA (0.85) imaging objective. (bottom left) A 3.5 mm diameter TiO2 metasurface in the Fourier plane generates a $600\times 600$ trap array in the imaging plane, spanning a 1.5 mm $\times$ 1.5 mm area, imaged on a CCD camera. b, Full image of the array, stitched together from 126 individual high-resolution images (see SI). Here, 8 high-resolution images of the edges are shown, illustrating the uniformity and quality of the traps in the array. c, Histogram of the optical power in individual tweezers. The power is determined by summing the intensity in a region around each trap center.
@@
In this section, we discuss the scalability of metasurface optical tweezer arrays and show how their performance compares to other pixel-based beam shaping devices, such as DMDs and liquid crystal SLMs. We find that small pixel sizes and a large pixel number, as provided by metasurfaces, are critically important to enable large and uniform tweezer arrays.
@@
To illustrate this point, we first consider reproducing a simple lens with a pixel-based device, as shown in Fig. 5a. The pixel size is denoted by $d$, the pixel density is $1/d$, and the wavelength of the incident laser beam is denoted by $\lambda$. We first convert a spherical lens to a Fresnel lens by taking the phase profile $\phi(x)$ of the former, modulo $2\pi$. When the phase profile of the Fresnel lens is approximated with a pixel-based device, the finite sampling constrains the steepness of phase gradients that can be reproduced, $\partial\phi/\partial x\lesssim\pi/2d$. This limits the ability to reproduce steep phase gradients at the edge of the lens. A larger pixel size will reduce the attainable phase gradient, reducing the usable diameter of the device, and effectively reducing the NA that can be achieved. Based on this argument, we derive an approximate analytical expression for the effective numerical aperture attainable for a given pixel-size-to-wavelength ratio, $\mathrm{NA}=1/\sqrt{(1+(4d/\lambda)^{2})}$ (see Methods for details). In Fig. 5b, we plot the effective NA for a broad range of pixel densities and several common tweezer wavelengths in the visible and near-infrared, illustrating that a smaller pixel size relative to the laser wavelength can lead to a dramatically increased effective NA. For a pixel size of several µm (typical for DMDs and liquid cyrstal SLMs), in the regime $d\gg\lambda$, the NA is limited to below 0.05. For subwavelength pixels, in the regime $d\lesssim\lambda$, NAs of 0.5 and higher can be reached.